OpenStudy (omgthatgirl):

what is the solution of the equation -3x^2+6x=-12 a. x=-1+sqrt5 b. x=1+sqrt5 c. x=-1+i sqrt5 d. x=1+i sqrt5

Nnesha (nnesha):

\[\huge\rm x = \frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a }\] quadratic formula first add 12 both side to write that equation in standard form Ax^2 + Bx+C=0 and substitute a ,b,c values into the formulathen solve :-)

OpenStudy (omgthatgirl):

ok..

Nnesha (nnesha):

a= ? b= ? c= ?

OpenStudy (omgthatgirl):

-3x^2 6x -12

OpenStudy (omgthatgirl):

?

Nnesha (nnesha):

coefficient at front of variable

Nnesha (nnesha):

\[\huge\rm \color{green}{A}x^2 +\color{green}{B}x +\color{green}{C}=0\] c is constant x^2 and x is variable a and b just numbers at front of variables

OpenStudy (omgthatgirl):

k

Nnesha (nnesha):

\[-3x^2 +6x = -12 \] first write it instandard form add 12 both side